Jose recently took a history test that had certain questions worth three points each. He correctly answered the same number of 3-point questions as 5-point questions, and earned a score of 72 points. How many 5-point questions did he answer correctly?
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

Jose recently took a history test that had certain questions worth three points each.
He correctly answered the same number of 3-point questions as 5-point questions, and earned a score of 72 points.
How many 5-point questions did he answer correctly?
:
let a = no. of 3 pt and no. of 5 pt questions
3a + 5a = 72
8a = 72
a = 72/8
a = 9 ea 5 pt questions (and 9 ea 3 point questions)
:
:
:
Check: 3(9) + 5(9) = 72

Answer by ikleyn(52794)

.
The condition to this problems looks more than strange.

It looks to be defective.